# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v3

@@ Abstract
Diffusion Models achieve state-of-the-art performance in generating new samples but lack a low-dimensional latent space that encodes the data into meaningful features. Inversion-based methods address this by reversing the denoising trajectory, mapping each image back to its approximated starting noise. In this work, we thoroughly analyze this procedure and focus on the relation between the initial Gaussian noise, the generated samples, and their corresponding latent encodings obtained through the DDIM inversion. First, we show that latents exhibit structural patterns in the form of less diverse noise predicted for smooth image regions. As a consequence of this divergence, we present that the space of image inversions is notably less manipulative than the original Gaussian noise. Next, we explain the origin of the phenomenon, demonstrating that, during the first inversion steps, the noise prediction error is much more significant for the plain areas than for the rest of the image. As a surprisingly simple solution, we propose to replace the first DDIM Inversion steps with a forward diffusion process, which successfully decorrelates latent encodings, leading to higher quality editions and interpolations. The code is available at https://github.com/luk-st/taba.
@@ 1 Introduction
Diffusion-based probabilistic models (DMs) [49], have achieved state-of-the-art results in many generative domains including image [8], speech [43], video [17], and music [34] synthesis. Nevertheless, one of the significant drawbacks that distinguishes diffusion-based approaches from other generative models like Variational Autoencoders (VAEs) [26] is the lack of an implicit latent space that encodes the images into low-dimensional, interpretable, or editable representations. To mitigate this issue, several works seek meaningful relations in the starting noise used for generations. This method, known as an inversion technique, was introduced with Denoising Diffusion Implicit Models (DDIM) [50], but led to the proliferation of works [12, 38, 19, 35, 18, 41]. The core idea is to use a denoising diffusion model to predict noise and add it to the image instead of subtracting it. Repeating this process effectively traces the backward diffusion trajectory, approximating the original noise that could have generated the image. However, due to approximation errors and biases introduced by the trained denoising model, discrepancies arise between the initial Gaussian noise and the reconstructed latent representation.
@@
While recent works [12, 38, 41, 19, 63] try to improve the inversion procedure from the perspective of tasks such as image reconstruction, editing, or interpolation, in this work, we focus on the inversion process itself and analyze the errors DDIM inversion introduces. To that end, we analyze the relation between sampled Gaussian noise, generated images, and their inverted latent encodings. First, we review existing studies and conduct additional analyses demonstrating that the reverse DDIM technique produces latent representations with pixel correlations that deviate significantly from a Normal distribution. As presented in fig. 1, we experimentally show that this deviation manifests as lower diversity in latent codes, particularly in regions corresponding to smooth image surfaces. We further attribute this discrepancy to the observation that the approximation error of DDIM inversion is higher for pixels representing smooth areas than for other parts of the image. Additionally, we show that this error discrepancy occurs predominantly during the initial steps of the inversion process. To present the consequences of the observed divergence, we analyze the effects of using latent variables instead of raw noise in practical DM applications. Our results show that the latent space is less adaptable for manipulation than the noise space. This limitation is particularly noticeable in lower-quality interpolations between randomly chosen latents and less expressive edits, especially in smooth regions of source images.
Finally, to mitigate the issue, we propose a surprisingly simple solution, which is replacing the first steps of the DDIM inversion process with a forward diffusion. In the final experiments, we show that such an approach successfully decorrelates the resulting latents, mitigating observed limitations, without degrading the reconstruction quality. Our main contributions can be summarized as follows:
@@
  • We show that reverse DDIM produces latent representations that deviate from the Normal distribution mostly because of less diverse values for the parts of the latents corresponding to the plain image surfaces.
  • We explain the origin of this phenomenon, showing that the inversion approximation error is more notable for the plain pixel areas, especially during the first noising steps.
  • We show that as a consequence of this divergence, the resulting latent space is less manipulative, which leads to the lower quality of interpolations and edits generated from latent variables as a starting state.
@@
  • We propose a simple yet effective solution to the problem by substituting early DDIM inversion steps with forward diffusion.
@@ 2 Background and Related Work
The training of DMs consists of forward and backward diffusion processes, where in the context of Denoising Diffusion Probabilistic Models (DDPMs [16]), the former one with training image $x_{0}$ and a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, can be expressed as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ with $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$.
@@
where $\epsilon^{(t)}_{\theta}(x_{t},c)$ is an output of a neural network (such as U-Net), and can be expressed as a combination of clean image ($x_{0}$) prediction, a direction pointing to previous denoising step ($x_{t}$), and a stochastic factor $\sigma_{t}z_{t}$, where $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$. In the standard DDPM model [16], the $\eta$ parameter is set to $\eta=1$. However, changing it to $\eta=0$ makes the whole process a deterministic Denoising Diffusion Implicit Model (DDIM) [50], a class of DMs we target in this work.
@@
$x_{t}={\sqrt{\alpha_{t}}}{x_{t-1}}+{(\sqrt{1-\bar{\alpha}_{t}}-{\sqrt{\alpha_{t}-\bar{\alpha}_{t}}})}\cdot{\epsilon^{(t)}_{\theta}(x_{t},c)}$ (2)
However, due to circular dependency on $\epsilon^{(t)}_{\theta}(x_{t},c)$, Dhariwal and Nichol [8] propose to approximate this equation by assuming the local linearity between directions $(x_{t-1}\rightarrow x_{t})$ and $(x_{t}\rightarrow x_{t+1})$, so that the model’s prediction in $t$-th inversion step can be approximated using $x_{t-1}$ as an input, i.e.,
@@
While such approximation is often sufficient to obtain good reconstructions of images, it introduces the error dependent on the difference $(x_{t}-x_{t-1})$, which can be detrimental for models that sample images with a few diffusion steps or use the classifier-free guidance [15, 38] for better prompt adherence. As a result, also noticed by recent works [12, 41], latents resulting from DDIM inversion do not follow the definition of Gaussian noise because of the existing correlations. In this work, we empirically study this phenomenon and explain its origin. We discuss the relations between the following three variables: Gaussian noise (${\mathbf{x_{T}}}$, an input to generate an image through a backward diffusion process), image sample (${\mathbf{x_{0}}}$, the outcome of the diffusion model generation process), and latent encoding (${\mathbf{\hat{x}_{T}}}$, the result of the DDIM inversion procedure as introduced in eq. 2).
@@
The DDIM inversion, despite the noise approximation errors, forms the foundation for many applications, including inpainting [58], interpolation [8, 63], and edition [53, 23, 14, 3, 7]. Several works [38, 12, 19, 35, 18, 37, 13, 4, 9, 62, 41, 54, 55, 40, 56, 2, 32, 20] aim to reverse the denoising process in text-to-image models, where prompt embeddings can strongly affect the final latent representation through Classifier-free-guidance (CFG) [15]. Null-text inversion [38] extends the DDIM inversion with additional null-embedding optimization, reducing the image reconstruction error. Other techniques improve inversion for image editing through seeking embeddings [37, 13, 9] or leveraging DDIM latents [4] for guidance. On the other hand, some works leverage additional numerical methods [35, 40, 12] to minimize inversion error. In particular, Renoise [12] iteratively improves the estimation of the next point along the forward diffusion trajectory by averaging multiple noise predictions with an additional patch-level regularization term penalizing pixel pairs’ correlations to ensure the latents’ editability. Huberman-Spiegelglas et al. [19] followed by Brack et al. [2] propose inversion methods for DDPMs, enabling the creation of various image edition results via inversion. Finally, to reduce the discrepancy between DDIM latents and Gaussian noises, Parmar et al. [41] propose to additionally regularize final DDIM Inversion outputs for better image editing, Lin et al. [32] introduce an alternative noise scheduler to improve inversion stability, while Hong et al. [18] propose an exact inversion procedure for higher-order DPM-Solvers, solving the optimization problem at each step.
3 Experiments
@@
Image res. $32\times 32$ $64\times 64$ $256\times 256$ $256\times 256$ $256\times 256$ $64\times 64$
Latent res. - - - $3\times 64\times 64$ $4\times 32\times 32$
@@
Table 1: Overview of diffusion models used for our experiments. We study both unconditioned and conditioned models, operating in pixel and latent spaces.
For thorough evaluation, we employ six different diffusion models working directly in the pixel space or latent space of the autoencoder described in table 1. For conditioning, we take prompts from the Recap-COCO-30K dataset [31] for IF and ImageNet class names for DiT. However, as noted by Mokady et al. [38], cfg introduces additional errors to the DDIM inversion, so to focus solely on the inversion approximation error in our setup, we disable cfg by setting the guidance scale to $w=1$. For both generation and inversion processes, we use the DDIM sampler with, unless stated otherwise, $T=100$ sampling steps. This choice [^1] We provide more discussion on the impact of sampling steps on observed inversion errors in appendix E. aligns with prior works [18, 12, 24] using from $50$ up to $200$ inversion steps as a proper balance between reconstruction quality and short algorithm runtime.
@@
Prior work and findings.
The reversed sampling (inversion) process provides the foundation for practical methods in many applications. The underlying assumption of those methods is that by encoding the image back with a denoising model, we can obtain the original noise that can be used for reconstruction. However, this assumption is not always fulfilled, which leads to the first question:
@@
Table 2: Mean of top-20 Pearson correlation coefficients inside $8\times 8$ patches for random Gaussian noises, latent encodings, and generations. Latents created with the reverse DDIM are much more correlated than noises.
@@
Figure 2: Latent encodings exhibit image patterns. For small pixel-space models (a), we observe correlations directly in the inversion results. For larger models (e.g., LDMs), the same patterns can be observed in the absolute errors between the latent and noise (b). This observation also holds for LDM models operating on $4$-channels, where we use PCA for visualization (c).
This question relates to several observations from the existing literature, which highlight that outputs of the DDIM inversion differ from the standard Gaussian noise [41, 12] and that the difference can be attributed to the approximation error [18, 55, 32]. In particular, Parmar et al. [41], followed by Garibi et al. [12], observe that latents “do not follow the statistical properties of uncorrelated, Gaussian white noise, causing poor editability”.
@@
First, we consolidate existing claims on the presence of correlations in the latent encodings (${\mathbf{\hat{x}_{T}}}$) by running an initial experiment that compares latents to images (${\mathbf{x_{0}}}$) and noises (${\mathbf{x_{T}}}$) across diverse diffusion architectures. In table 2, we calculate a mean of top-$20$ Pearson correlation coefficients (their absolute values) inside $C\times 8\times 8$ pixel patches, where $C$ is the number of channels (pixel RGB colors or latent space dimensions for latent models). The results, calculated with $10000$ samples, confirm that latent representations have significantly more correlated pixels than the independent Gaussian noises. In fig. 2, we show how the measured correlations manifest themselves in the latent codes calculated with inverse DDIM. For pixel models such as Deepfloyd IF, we observe clear groups of correlated pixels as presented in fig. 2. For latent diffusion models, we can highlight the inversion error by plotting the difference between the latent and the noise, as presented in fig. 2. This property also holds for LDMs with a $4$-channel latent space, where we use PCA to show the differences between the latents and noise for the DiT model (fig. 2).
@@ Conclusion.
Recent studies, supported by our initial experiments, demonstrate that latent representations computed using the DDIM inversion deviate from the expected characteristics of independent Gaussian noise. Specifically, both visual analysis and quantitative evaluations reveal significant correlations between the neighboring pixels.
@@ 3.2 Location of latent encodings space
In this work, we propose to delve further and empirically analyze the nature of this issue. To that end, we first pose the following question:
Research Question 2: Is there a common characteristic of the latent encodings that distinguishes them from Gaussian noise?
@@
To answer this question, we first investigate the geometric aspect of how the latent encodings are distributed in relation to the original noise and generated samples. In table 3, we determine the most probable angles formed by noises ${\mathbf{x_{T}}}$, samples ${\mathbf{x_{0}}}$, and latents ${\mathbf{\hat{x}_{T}}}$ for each model by binning the empirically observed angles and selecting those that satisfy the triangle angle sum, while maximizing the joint probability. We provide details on this methodology and visualizations in appendix F. Results of the experiment show that the angle located at the image and noise vertices (accordingly $\angle{\mathbf{x_{0}}}$ and $\angle{\mathbf{x_{T}}}$) are always acute and, in almost every case, the angle by the latent vertex ($\angle{\mathbf{\hat{x}_{T}}}$) is obtuse. This property implies that, due to approximation errors in the reverse DDIM process, latents reside in proximity to, but with a measurable offset, from the shortest-path trajectory between the noise distribution and the generated image.
Model $\angle{\mathbf{x_{0}}}$ $\angle{\mathbf{x_{T}}}$ $\angle{\mathbf{\hat{x}_{T}}}$
ADM-32 $29$ $28$ $123$
ADM-64 $11$ $60$ $109$
ADM-256 $24$ $73$ $83$
IF $34$ $40$ $106$
LDM $2$ $76$ $102$
DiT $4$ $66$ $110$
Table 3: Most probable angles next to the Gaussian noise ${\mathbf{x_{T}}}$, generations ${\mathbf{x_{0}}}$, and latents ${\mathbf{\hat{x}_{T}}}$ across different diffusion models.
Figure 3: $l_{2}$-distances between intermediate diffusion denoising steps $x_{t}$ and points on the linear interpolation path from noise ${\mathbf{x_{T}}}$ to the inverted latents ${\mathbf{\hat{x}_{T}}}$. The consecutive intermediate generation steps along the sampling trajectory consequently approach the latent.
@@
However, the trajectory defined by the backward diffusion process does not necessarily follow the linear path, as DMs learn a highly non-linear score function, causing the denoising path to follow the curved structure of the data manifold [51]. Therefore, to strengthen our analysis further, we investigate the location of the latent with respect to the generation trajectory. To that end, we analyze the distance between the following steps $\{x_{t}\}_{t=T\dots 1}$ of the backward diffusion process and intermediate points on the linear interpolation path between the noise and the associated latent. We present the results of this experiment in fig. 3, where each pixel, with coordinates $(t,\lambda)$, is colored according to the $l_{2}$ distance between the intermediate trajectory step $x_{t}$ and the corresponding interpolation step. This distance can be expressed as $\lVert(1-\lambda){\mathbf{x_{T}}}+\lambda{\mathbf{\hat{x}_{T}}}-x_{t}\rVert_{2}$.
ADM-$32$ ADM-$64$ ADM-$256$ IF LDM DiT Noise
$0.18_{\pm 0.33}$ $0.71_{\pm 1.05}$ $0.82_{\pm 1.27}$ $0.29_{\pm 0.53}$ $0.98_{\pm 1.40}$ $0.89_{\pm 1.29}$ $1.00$
Table 4: Variance of the latent encodings (${\mathbf{\hat{x}_{T}}}$) calculated with different models. In comparison, by definition, the variance of initial Gaussian noise equals $1.0$.
@@
We observe that, while moving from the random Gaussian noise (${\mathbf{x_{T}}}$) towards the final sample (${\mathbf{x_{0}}}$), the intermediate steps $x_{t}$ approach the latent (${\mathbf{\hat{x}_{T}}}$), while after the transition point present around 50-70% of the generative trajectory the distance to the latent becomes lower than the distance to the starting noise. This observation reveals that latent encodings retain some characteristics of the original input samples and contain information about the source generation.
To more closely localize the latent encodings, we propose to study their relations to the center of the analyzed space, point 00 being the mean of the initial noise ${\mathbf{x_{T}}}$. In table 4, we show the average variance of the latent encodings for randomly sampled and inverted 10,000 images. Latent encodings ${\mathbf{\hat{x}_{T}}}$ have, in general, lower variance than the noise, which means that the inversion process drives them towards the mean of the Gaussian distribution. The remaining question is, how does it manifest?
The visualizations shown in fig. 2 indicate that, in some cases, we can distinguish the structure of the original objects in the latent encodings. In particular, while the pixels associated with the objects have high diversity, which is also expected for white Gaussian noise, the areas related to the background are much smoother. This finding leads to the hypothesis that the most significant difference between the initial Gaussian noises and latent encodings is the limited variance in the areas corresponding to the background of the generated images. To validate it, we compare the properties of the resulting latents between plain and non-plain areas in the image. We determine binary masks $\mathcal{M}_{p}$
Model Mean Absolute Error Standard Deviation
@@
Table 5: Average per-pixel absolute error (between noise and latent) and standard deviation of the latents’ pixels corresponding to the plain and non-plain areas of generations. The error for plain pixels, where latents are less diverse, is higher than for other regions.
that select smooth surfaces in the image, according to the neighbor differences described in appendix G. In table 5, we show that the error between the starting noise ${\mathbf{x_{T}}}$ and the latent ${\mathbf{\hat{x}_{T}}}$ resulting from the DDIM inversion is higher for pixels corresponding to the plain areas. Across all the models, this trend goes along with a decrease in the standard deviation of the latents’ pixels related to those image regions. It suggests that DDIM inversion struggles with reversing the plain image areas, bringing them to 0 (average) and reducing their diversity.

### v4

@@ Abstract
Diffusion Models achieve state-of-the-art performance in generating new samples but lack a low-dimensional latent space that encodes the data into editable features. Inversion-based methods address this by reversing the denoising trajectory, transferring images to their approximated starting noise. In this work, we thoroughly analyze this procedure and focus on the relation between the initial noise, the generated samples, and their corresponding latent encodings obtained through the DDIM inversion. First, we show that latents exhibit structural patterns in the form of less diverse noise predicted for smooth image areas (e.g., plain sky). Through a series of analyses, we trace this issue to the first inversion steps, which fail to provide accurate and diverse noise. Consequently, the DDIM inversion space is notably less manipulative than the original noise. We show that prior inversion methods do not fully resolve this issue, but our simple fix, where we replace the first DDIM Inversion steps with a forward diffusion process, successfully decorrelates latent encodings and enables higher quality editions and interpolations. The code is available at https://github.com/luk-st/taba.
@@ 1 Introduction
Diffusion-based probabilistic models (DMs), (Sohl-Dickstein et al., 2015), have achieved state-of-the-art results in many generative domains including image (Dhariwal & Nichol, 2021), speech (Popov et al., 2021), video (Ho et al., 2022), and music (Liu et al., 2021) synthesis. Nevertheless, one of the significant drawbacks that distinguishes diffusion-based approaches from other generative models like Variational Autoencoders (Kingma & Welling, 2014) is the lack of an implicit latent space that encodes the images into low-dimensional, interpretable, or editable representations.
To mitigate this issue, several works seek meaningful relations in the approximated starting noise used for generations. This method, known as an inversion technique, was introduced by Song et al. (2021) with Denoising Diffusion Implicit Models (DDIM), and led to the proliferation of works (Garibi et al., 2024; Mokady et al., 2023; Huberman-Spiegelglas et al., 2024; Samuel et al., ; Hong et al., 2024; Parmar et al., 2023). The core idea is to use the noise predicted by the Diffusion Model and add it to the image instead of subtracting it. Repeating this process effectively traces the backward diffusion trajectory, approximating the noise that could have generated the image. However, due to approximation errors and biases introduced by the trained model, discrepancies arise between the original noise and its reconstruction – latent representation.
@@
While recent works (Garibi et al., 2024; Mokady et al., 2023; Parmar et al., 2023; Huberman-Spiegelglas et al., 2024; Zheng et al., 2024) try to improve the inversion procedure from the perspective of tasks such as image reconstruction, editing, or interpolation, in this work, we focus on the inversion process itself and analyze the errors DDIM inversion introduces. To that end, we analyze the relation between sampled noise, generated images, and their inverted latent encodings. First, we review existing studies and conduct additional analyses demonstrating that the reverse DDIM technique produces latent representations with pixel correlations that deviate significantly from a Normal distribution. As presented in Figure 1, we experimentally show that this deviation manifests as lower diversity in latents, particularly in regions corresponding to smooth image surfaces. We further attribute this discrepancy to the noise approximations in the first few inversion steps. We demonstrate that the inversion error is significantly higher and the predictions are notably less diverse for smooth image areas than for other regions.
To highlight the consequences of the observed divergence, we show that the DDIM-inversion-based latent space is less manipulative than the ground truth noise. This limitation is particularly noticeable in lower-quality image interpolations and less expressive edits, especially in smooth input image regions. Furthermore, we demonstrate that prior inversion methods, although designed to improve image reconstruction, fail to preserve the Gaussian properties of the latents. However, based on our analyses, we evaluate a simple fix, where we replace the first steps of the DDIM inversion process with a forward diffusion. In the final experiments, we show that such an approach successfully decorrelates the resulting latents, mitigating observed limitations without degrading the reconstruction quality. Our main contributions can be summarized as follows:
@@
  • We show that DDIM latents deviate from the Gaussian distribution, mostly because of less diverse noise predictions for the plain image surfaces during the first inversion steps.
  • We show that, consequently, the DDIM latents are less manipulative, leading to the lower quality of image interpolations and edits.
@@
  • We demonstrate that prior inversion methods do not address this issue and propose a simple and effective fix by substituting early inversion steps with a forward diffusion.
@@ 2 Background and Related Work
The training of DMs consists of forward and backward diffusion processes, where, in the context of Denoising Diffusion Probabilistic Models (DDPMs, Ho et al. (2020)), the former one with training image $x_{0}$ and a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, can be expressed as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ with $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$.
@@
where $\epsilon^{(t)}_{\theta}(x_{t},c)$ is an output of a neural network (such as U-Net), and can be expressed as a combination of clean image ($x_{0}$) prediction, a direction pointing to previous denoising step ($x_{t}$), and a stochastic factor $\sigma_{t}z_{t}$, where $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$. In the standard DDPM model, the $\eta$ parameter is set to $\eta=1$. However, changing it to $\eta=0$ makes the whole process a deterministic Denoising Diffusion Implicit Model (DDIM, Song et al. (2021)), a class of DMs we target in this work.
@@
$x_{t}={\sqrt{\alpha_{t}}}{x_{t-1}}+{(\sqrt{1-\bar{\alpha}_{t}}-{\sqrt{\alpha_{t}-\bar{\alpha}_{t}}})}\cdot{\epsilon^{(t)}_{\theta}(x_{t},c)}.$ (2)
However, due to circular dependency on $\epsilon^{(t)}_{\theta}(x_{t},c)$, Dhariwal & Nichol (2021) propose to approximate this equation by assuming the local linearity between directions $(x_{t-1}\rightarrow x_{t})$ and $(x_{t}\rightarrow x_{t+1})$, so that the model’s prediction in $t$-th inversion step can be approximated using $x_{t-1}$ as an input, i.e.,
@@
While such approximation is often sufficient to obtain good reconstructions of images, it introduces the error dependent on the difference $(x_{t}-x_{t-1})$, which can be detrimental for models that sample images with a few diffusion steps or use the classifier-free guidance (Ho & Salimans, 2021; Mokady et al., 2023) for better prompt adherence. As a result, also noticed by recent works (Garibi et al., 2024; Parmar et al., 2023), latents resulting from DDIM inversion do not follow the definition of Gaussian noise because of the existing correlations. In this work, we empirically study this phenomenon and explain its origin. We discuss the relations between the following three variables: Gaussian noise ($\mathbf{x_{T}}$}, an input to generate an image through a backward diffusion process), image sample ($\mathbf{x_{0}}$}, the outcome of the diffusion model generation process), and latent encoding ($\mathbf{\hat{x}_{T}}$}, the result of the DDIM inversion procedure as introduced in Eq. 2).
@@
The DDIM inversion, despite the noise approximation errors, forms the foundation for many applications, including inpainting (Zhang et al., 2023a), interpolation (Dhariwal & Nichol, 2021; Zheng et al., 2024), and edition (Su et al., 2022; Kim et al., 2022a; Hertz et al., 2022; Ceylan et al., 2023; Deja et al., 2023). Several works (Mokady et al., 2023; Garibi et al., 2024; Huberman-Spiegelglas et al., 2024; Samuel et al., ; Hong et al., 2024; Miyake et al., 2023; Han et al., 2024; Cho et al., 2024; Dong et al., 2023; Zhang et al., 2023b; Parmar et al., 2023; Tang et al., 2024; Wallace et al., 2023; Pan et al., 2023; Wang et al., 2024; Brack et al., 2024; Lin et al., 2024; Ju et al., 2024) aim to reverse the denoising process in text-to-image models, where prompt embeddings can strongly affect the final latent representation through Classifier-free-guidance (CFG) (Ho & Salimans, 2021). Null-text inversion (Mokady et al., 2023) extends the DDIM inversion with additional null-embedding optimization, reducing the image reconstruction error. Other techniques improve inversion for image editing through seeking embeddings (Miyake et al., 2023; Han et al., 2024; Dong et al., 2023) or leveraging DDIM latents (Cho et al., 2024) for guidance. On the other hand, some works leverage additional numerical methods (Samuel et al., ; Pan et al., 2023; Garibi et al., 2024) to minimize inversion error. In particular, Renoise (Garibi et al., 2024) iteratively improves the estimation of the next point along the forward diffusion trajectory by averaging multiple noise predictions, incorporating an additional patch-level regularization term that penalizes correlations between pixel pairs to ensure the editability of the latents. Huberman-Spiegelglas et al. (2024) followed by Brack et al. (2024) propose inversion methods for DDPMs, enabling the creation of various image edition results via inversion. Finally, to reduce the discrepancy between DDIM latents and Gaussian noises, Parmar et al. (2023) propose to additionally regularize final DDIM Inversion outputs for better image editing, Lin et al. (2024) introduce an alternative noise scheduler to improve inversion stability, while Hong et al. (2024) propose an exact inversion procedure for higher-order DPM-Solvers, solving the optimization problem at each step.
3 Analysis
@@
Image res. $32\times 32$ $64\times 64$ $\times 256 256×256$$\times 256$ $256\times 256$ $64\times 64$
Latent res. - - - $3\times 64\times 64$ $4\times 32\times 32$ -
@@
Table 1: Overview of diffusion models used for our experiments. We study both unconditioned and conditioned models, operating in pixel and latent spaces. More details on models are provided in Appendix G.
In our experiments, we employ six different diffusion models, which we compare in Table 1. For both generation and inversion processes, we use the DDIM sampler with, unless stated otherwise, $T=100$ steps. We provide more details on the number of diffusion steps in Appendix F.
@@
The inversion process provides the foundation for practical methods in many applications, with the underlying assumption that by encoding the image back with a denoising model, we can obtain the original noise that can be used for reconstruction. However, this assumption is not always fulfilled, which leads to our first question:
@@
Table 2: Mean of top-20 Pearson correlation coefficients inside $8\times 8$ patches for random Gaussian noises, latent encodings, and generations. DDIM Latents are much more correlated than noises.
@@
Figure 2: Latent encodings exhibit image patterns. For small pixel-space models (a), we observe correlations directly in the inversion results. For larger models (e.g., LDMs), the same patterns can be observed in the absolute errors between the latent and noise (b). This observation also holds for LDM models operating on 44-channels, where we use PCA for visualization (c).
Prior work and findings.
This question relates to several observations from the existing literature, which highlight that outputs of the DDIM inversion differ from the standard Gaussian noise (Parmar et al., 2023; Garibi et al., 2024) and that the difference can be attributed to the approximation error (Hong et al., 2024; Wallace et al., 2023). While these works notice the divergence between noise and latent encodings, they do not validate them or study the source of this issue.
@@
First, we consolidate existing observations on the presence of correlations in the latent encodings ($\mathbf{\hat{x}_{T}}$}) and validate them by running an initial experiment that compares latents to images ($\mathbf{x_{0}}$}) and noises ($\mathbf{x_{T}}$}) across diverse diffusion architectures. In Table 2, we calculate a mean of top-$20$ Pearson correlation coefficients (their absolute values) inside $C\times 8\times 8$ pixel patches, where $C$ is the number of channels (pixel RGB colors or latent space dimensions for latent models). The results confirm that latent representations have significantly more correlated pixels than the noise. In Fig. 2, we show how the measured correlations visually manifest themselves in the latents. For pixel models such as Deepfloyd IF, we observe clear groups of correlated pixels as presented in Fig. 2a. For latent diffusion models, we can highlight the inversion error by plotting the difference between the latent and the noise, as presented in Fig. 2b. This property also holds for LDMs with a $4$-channel latent space, with the use of PCA (Fig. 2c).
@@ Conclusion.
Our initial experiments numerically validate observations from recent studies and demonstrate that latent representations computed using the DDIM inversion deviate from the expected characteristics of independent Gaussian noise. Specifically, both visual analysis and quantitative evaluations reveal significant correlations between the neighboring pixels.
@@ 3.2 Location of latent encodings space
To delve deeper, we first propose to empirically analyze the nature of this issue, posing a question:
Research Question 2: How do DDIM inversion latents differ from the Gaussian noise?
Figure 3: $l_{2}$-distances between intermediate denoising steps $x_{t}$ and points on the linear interpolation path from noise $\mathbf{x_{T}}$} to the inverted latents $\mathbf{\hat{x}_{T}}$}. The consecutive intermediate generation steps along the sampling trajectory consequently approach the latent.
@@
To answer this question, we first geometrically investigate the location of the latents with respect to the generation trajectory. To that end, we analyze the distance between the following steps $\{x_{t}\}_{t=T\dots 1}$ of the backward diffusion process and intermediate points on the linear interpolation path between the noise and the DDIM latent. We present the results of this experiment in Fig. 3, where each pixel, with coordinates $(t,\lambda)$, is colored according to the $l_{2}$ distance between the intermediate trajectory step $x_{t}$ and the corresponding interpolation step. This distance can be expressed as $\lVert(1-\lambda){\mathbf{x_{T}}}+\lambda{\mathbf{\hat{x}_{T}}}-x_{t}\rVert_{2}$.
@@
We observe that, while moving from the initial noise ($\mathbf{x_{T}}$}) towards the final sample ($\mathbf{x_{0}}$}), the intermediate steps $x_{t}$ approach the DDIM inversion latent ($\mathbf{\hat{x}_{T}}$}), while after the transition point around 50-70% (timesteps 50-30) of the generative trajectory, the distance to the latent becomes lower than the distance to the starting noise. This observation reveals that latents retain some characteristics of the original samples and contain information about the source generation.
Model Absolute Error Standard Deviation
@@
Table 3: Average per-pixel absolute error (between noise and latent) and standard deviation of the latents’ pixels corresponding to the plain and non-plain image areas. The error for plain pixels, where latents are less diverse, is higher than for other regions.
Similar observation can be made on the basis of visualizations in Fig. 2, where we can distinguish the coarse shape of the original objects in the latents. In particular, while the pixels associated with the objects have high diversity, the areas related to the background are much smoother. This leads to the hypothesis that the most significant difference between the initial noises and latent encodings is the limited variance in the areas corresponding to the background of the generated images. To validate it, we compare the properties of the latents between plain and non-plain areas in the image. We determine binary masks $\mathcal{M}_{p}$ that select smooth surfaces in the image, such as sky, sea, plain backgrounds, or surfaces (see Fig. 12 for examples), according to the neighbor differences as described in Appendix I. In Table 3, we show that the error between the starting noise $\mathbf{x_{T}}$} and the latent $\mathbf{\hat{x}_{T}}$} resulting from the DDIM inversion is higher for pixels corresponding to the plain areas. Across the models, this trend goes along with a decrease in the standard deviation of the latents’ pixels related to those regions. It suggests that DDIM inversion struggles with reversing the plain image areas, bringing them to mean (0) and reducing their diversity.